Julio's grandfather is 3 times his age. the total sum of their ages is 92. how old is Julio's grandfather?
Answer by stanbon(75887)

Julio's grandfather is 3 times his age. the total sum of their ages is 92. how old is Julio's grandfather?
----------
Equations:
g = 3j
g + j = 92
-------
Substitute for "g" and solve for J:
3j + j = 92
4j = 92
julio = 23 yrs
---
Solve for "g":
g = 3j
g = 3*23 = 69 years
==========================
Cheers,
Stan H.
==========================